# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ 9.3 Inconsistent surface registration
In this section, we test our SBN-Opt on the problem of inconsistent surface registration. In the application of inconsistent surface registration, our framework demonstrates consistently superior performance when compared to the landmark-dependent LSQC method proposed by Qiu et al qiu2020inconsistent. Across three distinct registration tasks, our method achieves mappings with lower angular distortion and improved feature alignment. For details of the problem formulation, see Sec.7.
The first example is the registration of partial tooth surfaces. In Fig.13, the first two columns show the conformal parameterization of the moving surface and the target static tooth surface, respectively. There are 4 regions selected on each surface for alignment and are highlighted in Fig.13. In this example, the weights of intensity matching loss $E_{\text{I}}$, landmark matching loss $E_{\text{pc}}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 1e-1, 5e-1 and 1e-3, respectively. We compare our result with that of the method in Qiuqiu2020inconsistent which is denoted by LSQC in the following. And to perform Qiu’s method, we select four landmarks, one in each region for two surfaces, respectively. As shown in Fig.14, When registering two partial tooth surfaces using four selected feature regions, our method reduced the angular distortion, measured by the average L2 norm of the BCs, by 11.3% (from 0.115 to 0.102). More significantly, the registration accuracy was substantially improved, with the average L1 intensity mismatch error on the overlapping region dropping by 16.9% (from 0.0794 to 0.0660). The registration result in the 2D domain is shown in the third column of Fig.13. The green mesh is the deformed mesh from the source mesh under the bijective deformation map with the blue highlighted regions overlapping orange highlighted regions in the target mesh. The intersection region of the two meshes is the region of correspondence amongst the two tooth surfaces, see Fig.15.
@@
Figure 14: Statistical comparison of registration results for the tooth surfaces (Fig.13, Fig.15). This figure format is used for all subsequent comparisons (Figs. 17 and 20). The histograms on the left (a, b) compare the angular distortion, measured by the magnitude of the BCs. The histograms on the right (c, d) compare the registration accuracy, measured by the percentage of intensity mismatch error in the overlapping region. In each pair, the result of SBN-Opt is on the left (a, c) and that of the LSQC method is on the right (b, d).
@@
The second example is the registration of two human faces, which are obtained from the FaMoS DatasetTMPEH:CVPR:2023. In Fig.16, the first two columns show the conformal parameterization of the moving surface and the target static tooth surface, respectively. There are 4 regions (two eyes, nose and mouth) selected on each surface for alignment and are highlighted in Fig.16. And to perform Qiu’s method, we select four landmarks, one in each region for two surfaces, respectively. In this example, the weights of intensity matching loss $E_{\text{I}}$, landmark matching loss $E_{\text{pc}}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 1e-1, 5e-1 and 1e-3, respectively. We compare our result with that of LSQC. As shown in Fig.17, our method yielded even more pronounced improvements in geometric quality. The average L2 norm of the BCs was reduced by 35.7% (from 0.111 to 0.0711), indicating a much smoother mapping. This was accompanied by a 10.7% decrease in the intensity mismatch error on the overlapping region of correspondence (from 0.0242 to 0.0216). The registration result in the 2D domain is shown in the third column of Fig.16. The green mesh is the deformed mesh from the source mesh under the bijective deformation map with the blue highlighted regions overlapping orange highlighted regions in the target mesh. The intersection region of the two meshes is the region of correspondence amongst the two tooth surfaces, see Fig.18.
Figure 16: The first two columns show the conformal parametrizations of the moving and static human faces, respectively. The region correspondences in the 2D parameter domains are also displayed. The registration result in the 2D domain is shown in the rightmost column.
Figure 17: Quantitative comparison for the full human face registration shown in Fig.16. The meaning of each histogram corresponds to that in Fig.14.
Figure 18: Visualization of the computed correspondence for the full human face registration.
The third case is an example of partial surface registration. Two human faces are cut into two partial faces, both of whose nose and mouth are remained. In Fig.19, the first two columns show the conformal parameterization of the moving surface and the target static tooth surface, respectively. There are 3 regions (bridge of nose, nose and mouth) selected on each surface for alignment and are highlighted in Fig.19. And to perform Qiu’s method, we select seven landmarks, two in bridge of nose, three in nose and two in mouth, respectively. In this example, the weights of intensity matching loss $E_{\text{I}}$, landmark matching loss $E_{\text{pc}}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ are 1, 2e-1, 1e-3 and 1e-3, respectively. We compare our result with that of LSQC. As shown in Fig.20, our approach demonstrated its most significant advantage in reducing distortion. The angular distortion was reduced by an impressive 39.4% (from 0.1464 to 0.0887). The average intensity mismatch error was also reduced by 3.06% (from 0.0327 to 0.0317). Qualitatively, our method produced a more plausible and reasonable correspondence; the resulting overlap region in the source mesh correctly avoided the eye region, an anatomical feature not present in the target partial face. In contrast, the LSQC method produced an overlap that incorrectly extended to the eye region, see Fig.21. The registration result in the 2D domain is shown in the third column of Fig.19. The green mesh is the deformed mesh from the source mesh under the bijective deformation map with the blue highlighted regions overlapping orange highlighted regions in the target mesh.
@@
Figure 20: Quantitative comparison for the partial human face registration shown in Fig.19 The meaning of each histogram corresponds to that in Fig.14.
Figure 21: Visualization of the computed correspondence for the partial human face registration. The identified region of correspondence by our method is highlighted in blue on both the source surface (left) and the target surface (center). The complete surfaces are shown in yellow. Compared with the overlap identified by LSQC on the source mesh (right), computed overlap by our method correctly avoids mapping the eye region from the source, which is not present in the target.
@@ 10 Conclusions
In this work, we introduced SBN-Opt, a novel neural-based framework for free-boundary diffeomorphism optimization, and its core engine, the Spectral Beltrami Network (SBN). By rigorously establishing the mathematical properties of the Least Squares Quasiconformal (LSQC) energy including existence, uniqueness, similarity-invariance, and resolution-independence, we laid a solid theoretical foundation for SBN-Opt. SBN is a neural surrogate that, for the first time, faithfully approximates the free-boundary LSQC solver. Its unique architecture, which synergistically combines multiscale message-passing with mesh spectral layers, overcomes the limitations of prior methods by effectively capturing both local and global geometric dependencies.
The SBN-Opt framework leverages this powerful surrogate to reframe the complex task of diffeomorphism optimization into a gradient-based problem over the space of BCs and pinned-point conditions. This paradigm shift provides an unprecedented level of explicit control over local geometric distortion while elegantly handling the challenges of free-boundary conditions without the need for artificial constraints or landmark-based scaffolding.
Extensive experiments on density equalization and inconsistent surface registration demonstrate the clear superiority of our approach. SBN-Opt consistently produces mappings with significantly lower distortion and higher accuracy compared to established numerical algorithms. By bridging the gap between the mathematical rigor of quasiconformal geometry and the computational flexibility of deep learning, SBN-Opt provides a powerful, theoretically-grounded, and versatile tool. This work not only addresses long-standing challenges in surface mapping but also paves the way for future advancements in geometry processing, medical imaging, and scientific visualization.

### v2

@@ 8.3 Inconsistent surface registration
In this section, we tested our SBN-Opt on the problem of inconsistent surface registration. In the application of inconsistent surface registration, our framework demonstrates consistently superior performance when compared to the landmark-dependent LSQC method proposed by Qiu et al.Qiu and Lui (2020). Across three distinct registration tasks, our method achieves mappings with lower angular distortion and comparative or slightly better feature alignment. For details of the problem formulation, see Section 6.
The first example is the registration of tooth surfaces. In Figure 16, the first two columns show the conformal parameterizations of the moving surface and the target static tooth surface, respectively. There were 4 regions selected on each surface for alignment and are highlighted in Figure 16. In this example, the weights of intensity matching loss $E_{\text{I}}$, landmark matching loss $E_{\text{pc}}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ were 1, 1e-1, 5e-1 and 1e-3, respectively. We compared our result with that of the method in QiuQiu and Lui (2020) which is denoted by LSQC in the following. And to perform Qiu’s method, we selected four landmarks, one in each selected region on both surfaces. As shown in Figure 17, When registering two tooth surfaces using four selected feature regions, our method reduced the angular distortion, measured by the average $\ell^{2}$ norm of BCs, by 12.2% (from 0.1135 to 0.09966). Additionally, the angle distortion histogram for SBN-Opt was more concentrated around lower values and exhibits a lighter tail than that of LSQC, with mean angle distortion reduced by 10.7% ($6.943^{\circ}$ to $6.202^{\circ}$). More significantly, the registration accuracy was substantially improved, with the average L1 intensity mismatch error on the overlapping region dropping by 16.9% (from 0.0794 to 0.0660). The registration result in the 2D domain is shown in the third column of Figure 16. The green mesh is the deformed mesh from the source mesh under the bijective deformation map with the blue highlighted regions overlapping orange highlighted regions in the target mesh. The intersection region of the two meshes is the region of correspondence amongst the two tooth surfaces, see Figure 18.
@@
Figure 17: Statistical comparison of registration results for the tooth surfaces (Figure 16, Figure 18). This figure format is used for all subsequent comparisons (Figure 20 and Figure 23). Top row: results for LSQC; bottom row: results for SBN-Opt. From left to right: (a) Histogram of BC magnitudes; (b) Histogram of per-face angle distortion (in degrees); (c) Histogram of the intensity mismatch error distribution.
@@
The second example is the registration of two partial human faces, which are obtained from the FaMoS DatasetBolkart et al. (2023). Each subject was cropped so that the nose and mouth regions remain while the eyes and surrounding regions were partially removed. In Figure 19, the first two columns show the conformal parameterizations of the moving and target partial faces, respectively. There were 3 regions (bridge of nose, nose and mouth) selected on each surface for alignment and are highlighted in Figure 19. And to perform Qiu’s method, we selected seven landmarks, two in bridge of nose, three in nose and two in mouth, respectively. In this example, the weights of intensity matching loss $E_{\text{I}}$, landmark matching loss $E_{\text{pc}}$, angle distortion loss $E_{\text{BC}}$ and smoothness loss $E_{\text{smooth}}$ were 1, 5e-1, 1e-2 and 1e-3, respectively. We compared our result with that of LSQC. As shown in Figure 20, our approach demonstrates its most significant advantage in reducing distortion. The average BC norm was reduced by an impressive 23.7% (from 0.1439 to 0.1098). Consistent with this, the mean angle distortion was also decreased by 23.5% (from $9.123^{\circ}$ to $6.983^{\circ}$) The average intensity mismatch error was also reduced by 26.0% (from 0.0327 to 0.0242). Qualitatively, our method produced a more reasonable correspondence; the resulting overlap region in the source mesh correctly avoided the eye region, an anatomical feature not present in the target partial face. In contrast, the LSQC method produced an overlap that incorrectly extended to the eye region, see Figure 21. The registration result in the 2D domain is shown in the rightmost of Figure 19. The green mesh is the deformed mesh from the source mesh under the bijective deformation map with the blue highlighted regions overlapping orange highlighted regions in the target mesh.
@@
Figure 20: Quantitative comparison for the partial human face registration shown in Figure 19 The meaning of each histogram corresponds to that in Figure 17.
Figure 21: Overlap region comparison for partial human face registration with mouths closed. From left to right: overlap region visualized on the source mesh by SBN-Opt, on the target mesh by SBN-Opt, on the source mesh by LSQC, and on the target mesh by LSQC. Same Format is used in Figure 24. In all subfigures, the full facial surface is shown in yellow, while the overlap region identified by the registration is rendered in blue. Compared with the LSQC’s result, the overlap computed by SBN-Opt is more anatomically plausible, which correctly excludes the eye region on the source mesh and yields a more localized correspondence around the nose and mouth on the target mesh.
The third example further tested our method on partial human faces whose mouths are open, so that both the source and target surfaces are multiply connected. In this setting, the parameter domains contained interior boundary components corresponding to the mouth openings. Figure 22 shows the conformal parameterizations of the moving and static open-mouth faces in the first two columns, with highlighted feature regions used to guide registration, and the resulting 2D registration in the rightmost column. A key advantage of SBN-Opt is that it can operate directly on these multiply connected parameter domains without any topological modification. By contrast, the LSQC method of Qiu Qiu and Lui (2020) assumes simply connected domains; to apply it here, one must first manually fill the mouth openings on both surfaces, convert them into simply connected surfaces, and then register the filled meshes. We used the same loss components $E_{\text{I}}$, $E_{\text{pc}}$, $E_{\text{BC}}$ and $E_{\text{smooth}}$ as in the previous examples, with weights 1,5e-1,3e-1 and 0, respectively. Figure 23 reports the quantitative comparison between SBN-Opt and LSQC. SBN-Opt again yielded a more concentrated BC histogram with the mean value dropped by 37.1% (from 0.1078 to 0.0678), and the angle distortion was more concentrated around lower values with the mean value reduced by 30.8% (from $7.465^{\circ}$ to $5.169^{\circ}$). Meanwhile, the intensity matching result is comparable compared with that of LSQC. The corresponding 3D overlap regions are visualized in Figure 24. On both the source and target meshes, the overlap computed by SBN-Opt adhered closely to the facial features and correctly tracked the mouth opening boundaries. In contrast, the LSQC-based registration, which is working on filled, simply connected surrogates of the surfaces, exhibits less faithful alignment around the bridge of nose region and the overlap cannot correctly avoid the eye region in the target mesh. This example highlights that SBN-Opt naturally extends to multiply connected inconsistent surface registration without requiring domain filling or topological surgery, while simultaneously achieving better distortion control, competitive or superior matching accuracy and more reasonable registration result.
Figure 23: Quantitative comparison for the partial human face registration shown in Figure 19. The meaning of each histogram corresponds to that in Figure 17.
Figure 24: Visualization of the computed correspondence for the partial human face registration with mouths open. The identified region of correspondence by our method is highlighted in blue on both the source surface (left) and the target surface (center). The complete surfaces are shown in yellow. Compared with the overlap identified by LSQC on the target mesh (right), computed overlap by our method correctly avoids mapping the eye region from the source, which is not present in the target.
@@ 9 Conclusions
In this work, we introduced SBN-Opt, a neural framework for free-boundary diffeomorphism optimization built upon the Spectral Beltrami Network (SBN). By rigorously establishing the mathematical properties of the Least-squares Quasiconformal (LSQC) energy including existence and uniqueness of minimizers, similarity-invariance, and resolution independence, we laid a solid theoretical foundation for using LSQC in free-boundary settings. Additionally, we analyzed its stability and differentiability, which enables gradient-based optimization via LSQC. Due to the heavy computation cost, we propose the Spectral Beltrami Network. SBN is a neural surrogate that, for the first time, faithfully approximates the free-boundary LSQC solver. Its unique architecture, which synergistically combines multiscale message-passing with mesh spectral layers, overcomes the limitations of prior methods by effectively capturing both local and global geometric dependencies.
A distinctive advantage of our approach is that the two‐pinned‐point condition is treated as part of the input to SBN itself; consequently, both the Beltrami coefficients and the pinned points are optimization variables and the corresponding constraints remain differentiable. This contrasts with many existing quasiconformal geometry based and other diffeomorphic registration methods, which require externally prescribed landmark constraints in order to obtain nontrivial solutions, and thus cannot directly handle general free-boundary problems in a gradient-based pipeline.
Leveraging this design, SBN-Opt reframes the free-boundary diffeomorphism problem as an optimization over the space of admissible Beltrami coefficients and pinned-point conditions, enabling explicit control of local geometric distortion and boundary shapes without artificial scaffolds or fixed landmarks.
Extensive experiments on equiareal parameterization and inconsistent surface registration demonstrate the clear superiority of our approach. SBN-Opt consistently produces mappings with significantly lower distortion and higher accuracy compared to established numerical algorithms. By bridging the gap between the mathematical rigor of quasiconformal geometry and the computational flexibility of deep learning, SBN-Opt provides a powerful and theoretically-grounded tool. This work not only addresses long-standing challenges in surface mapping but also paves the way for future advances in geometry processing, medical imaging, and scientific visualization.